# Universal rapid machine learning models for predicting unconvoluted and convoluted X-ray Absorption Spectra

### v1

@@ 3 Results &\& Discussion
In practice, an XANES prediction model was constructed for individual elements, exemplified here using the K-edge XANES of Ni, S and Ru. Taking Ni as a representative case to evaluate model performance, three GNN architectures XAS3D, GraphNet[24], and SGN[25] were constructed for XANES prediction. To prevent performance degradation of the models due to improper hyperparameter selection, 72 sets of hyperparameters were searched for each model. For further details, see Support information section "Hyperparameters optimization". Specifically, three shared hyperparameters across all three models the number of layers in the model, the hidden channels per GNN layer, and the learning rate were sampled using an identical grid of 72 points. Meanwhile, other model-specific hyperparameters for each model were sampled via a hypercube sampling method. Fig.1 illustrates the distribution of MAEs across these configurations. Among the 72 sets of hyperparameters evaluated, XAS3D demonstrated a median value of 0.0169, outperforming both GraphNet (0.0578) and SGN (0.0433), indicating its overall superior performance. Under their respective optimal hyperparameter settings, the models achieved mean absolute errors (MAE) as follows: 0.0128 for XAS3D, 0.0389 for GraphNet, and 0.0397 for SGN. To comprehensively visualize performance variability across models, we ranked them according to their validation set metrics and plotted spectral predictions covering five quantiles (q10–q90), as illustrated in Fig.1. For Ni/XAS3D model: predictions showed visually indistinguishable agreement with theoretical spectra up to q70; subtle deviations emerged only at q90. In contrast, both GraphNet and SGN exhibited discernible discrepancies even at q10, with progressive divergence through q30–q70, reaching substantial mismatches in spectral trends by q90. The S/XAS3D model registered an MAE of 0.0197, higher than Ni/XAS3D model 0.0130, likely due to inherent challenges in capturing electronic structure effects without explicit modeling. As a soft X-ray probe (absorption edge at 2472 eV), sulfur XAS contains rich electronic information derivable solely from geometric inputs. In terms of spectral prediction performance, excellent agreement between predicted and theoretical spectra was observed up to the q50 quantile. However, the predictive accuracy began to deteriorate starting from the q70 quantile. At the q90 quantile, the predicted position of the white line peak within the range of 2475–2485 eV exhibits an energy shift towards higher values by 2.3 eV. For ruthenium representative of 4d transition metals or platinum-group elements, the Ru/XAS3D model achieved an exceptionally low MAE of 0.0082. Even at q90, predicted spectra aligned closely with theoretical one. Considering the core-hole width of the K-edge of Ru (5.33 eV) is significantly larger than that of Ni (1.44 eV) and S (0.59 eV). Besides the model’s favorable performance, this result is also attributable to the significant broadening of Ru element core-hole excitations, which obscures fine structural details of cross-section in theoretical XANES.

### v2

@@ 2 Method
3 Hyperparameters optimization of machine learning models
To prevent performance degradation of the models due to improper hyperparameter selection, 72 sets of hyperparameters were searched for each model. Specifically, a grid search was performed on three shared hyperparameters across all three GNN models: the number of layers (nl), the hidden embedding size (hidden) per GNN layer, and the learning rate (lr). Six values were tested for the number of layers (1, 2, 3, 4, 6, 8), six for the hidden embedding size (16, 32, 64, 96, 128, 256), and two for the learning rate (0.001, 0.0005). Through grid sampling across these ranges, a total of 72 paramter configurations were generated (6 × 6 × 2). Each GNN model’s unique hyperparameters were sampled in 72 sets, which were then combined with the shared hyperparameters to form 72 complete hyperparameter configurations for the given model. For instance, in the SGN model: The K value hyperparameter adopted nine discrete integers from 2 to 10 inclusive. By repeating this sequence eight times and applying random permutation, we obtained 72 distinct K value instances. When integrated with previously sampled general GNN hyperparameters, these yielded 72 comprehensive hyperparameter configurations for SGN. Analogously, the XAS3D model’s unique hyperparameters—including Distance embedding dim, Angle embedding dim, Middle embedding size, and Output block layer count—underwent independent sampling to produce 72 variants each. Integration with the with previously sampled general GNN hyperparameters resulted in 72 comprehensive hyperparameter configurations for XAS3D. We performed hyperparameter optimization for the XAS3D, GraphNet, and SGN models on the Ni XANES prediction dataset. The performance across 72 hyperparameter sets for these three models is presented in Fig.1. The performance boxplots across 72 hyperparameter sets for these three models are shown in Fig.1. Among the 72 sets of hyperparameters evaluated, XAS3D demonstrated a median value of 0.0169, outperforming both GraphNet (0.0578) and SGN (0.0433), indicating its overall superior performance. Under their respective optimal hyperparameter configurations, the models attained the following MAE scores on the testing set: 0.0137 (XAS3D), 0.0388 (GraphNet), and 0.0385 (SGN). On the validation set, the corresponding MAE metrics were 0.0128 (XAS3D), 0.0389 (GraphNet) and 0.0397 (SGN).
Figure 1: Boxplots of performance metric (MAE) of XAS3D, GraphNet and SGN GNN models using different hyperparameters in Ni K-edge XANES prediction.
@@ 4 Results &\& Discussion
In practice, an XANES prediction model was constructed for individual elements, exemplified here using the K-edge XANES of Ni, S and Ru. Taking Ni as a representative case to evaluate model performance, three GNN architectures XAS3D, GraphNet[24], and SGN[25] were constructed for XANES prediction. To prevent performance degradation of the models due to improper hyperparameter selection, 72 sets of hyperparameters were searched for each model. For further details, see Support information section "Hyperparameters optimization". Specifically, three shared hyperparameters across all three models the number of layers in the model, the hidden channels per GNN layer, and the learning rate were sampled using an identical grid of 72 points. Meanwhile, other model-specific hyperparameters for each model were sampled via a hypercube sampling method. Fig.2 illustrates the distribution of MAEs across these configurations. Among the 72 sets of hyperparameters evaluated, XAS3D demonstrated a median value of 0.0169, outperforming both GraphNet (0.0578) and SGN (0.0433), indicating its overall superior performance. Under their respective optimal hyperparameter settings, the models achieved mean absolute errors (MAE) as follows: 0.0128 for XAS3D, 0.0389 for GraphNet, and 0.0397 for SGN. To comprehensively visualize performance variability across models, we ranked them according to their validation set metrics and plotted spectral predictions covering five quantiles (q10–q90), as illustrated in Fig.2. For Ni/XAS3D model: predictions showed visually indistinguishable agreement with theoretical spectra up to q70; subtle deviations emerged only at q90. In contrast, both GraphNet and SGN exhibited discernible discrepancies even at q10, with progressive divergence through q30–q70, reaching substantial mismatches in spectral trends by q90. The S/XAS3D model registered an MAE of 0.0197, higher than Ni/XAS3D model 0.0130, likely due to inherent challenges in capturing electronic structure effects without explicit modeling. As a soft X-ray probe (absorption edge at 2472 eV), sulfur XAS contains rich electronic information derivable solely from geometric inputs. In terms of spectral prediction performance, excellent agreement between predicted and theoretical spectra was observed up to the q50 quantile. However, the predictive accuracy began to deteriorate starting from the q70 quantile. At the q90 quantile, the predicted position of the white line peak within the range of 2475–2485 eV exhibits an energy shift towards higher values by 2.3 eV. For ruthenium representative of 4d transition metals or platinum-group elements, the Ru/XAS3D model achieved an exceptionally low MAE of 0.0082. Even at q90, predicted spectra aligned closely with theoretical one. Considering the core-hole width of the K-edge of Ru (5.33 eV) is significantly larger than that of Ni (1.44 eV) and S (0.59 eV). Besides the model’s favorable performance, this result is also attributable to the significant broadening of Ru element core-hole excitations, which obscures fine structural details of cross-section in theoretical XANES. Building upon the good performance of single-element models, we conducted studies on multi-element models. Motivated by the above single-element performance, we next asked whether multi-element XANES prediction models possess potential advantages beyond their inherent generalization capability. Our findings demonstrate that multi-element models utilize knowledge acquired from other elements to improve the predictive performance for elements with limited data availability.
In practical applications, not every element has a large number of samples for the given absorption edge XAS to conduct independent model training. To simulate this scenario, we conducted a test with Ni as the research element, including 50-8000 samples. For comparison, baseline models were trained solely on the Ni element, with both the training and validation sets comprising only Ni element. To investigate the influence of samples from other elements in the training set during multi-element XAS prediction model training, besides Ni, 34,000 samples of four other 3d or 4d transition metal elements were included for model training. The validation set contained no Ni samples, consisting only of samples from other 3d or 4d transition metal elements. This design was implemented because, in practice, when samples of a specific element (e.g., Ni) are scarce, it is desirable to include as many of its samples as possible in the training set, leaving insufficient samples for the validation set. Consequently, this study evaluated the predictive performance of models optimized using non-Ni element validation sets on Ni XANES spectra, results are shown in Tab.1. It can be observed that when the dataset includes 34,000 samples of Fe, Co, Cu, and Zn elements, even with only 50 Ni samples included, the model can achieve a satisfactory prediction performance with an MAE of 0.0175. It is evident that samples from other 3d transition metal elements significantly enhance the prediction accuracy for Ni XANES spectra. Moreover, the enhancement ratio increased as the number of available Ni samples decreases. For instance, with only 50 Ni samples, incorporating 34,000 samples each of Fe, Co, Cu, and Zn into the training set reduced the model’s Mean Absolute Error (MAE) by 80% compared to training solely on Ni samples. This MAE reduction percentage decreased to 30 percentage points when the Ni sample count was 500, and further to 14% at counts of 2,000 or 4,000. When the Ni sample size was sufficiently large (8,000), adding extra 3d transition metal samples provided no additional improvement in model performance. In contrast, incorporating 4d transition metal samples showed negligible impact on Ni XANES prediction performance, except under conditions of extreme Ni scarcity (e.g., 50 samples).
Table 1: Performance metrics MAE in the testing dataset for XANES prediction of Ni element. MAE 3d/ MAE 4d denotes that the model training set comprises samples of Ni element and 34,000 samples of 3d/4d transition metals. MAE Ni indicates that the training set contains only Ni element samples.
Number of Ni samples
in training set MAE 3d MAE 4d MAE Ni
50 0.0175 0.0301 0.0869
500 0.0156 0.0244 0.0223
2000 0.0145 0.0165 0.0169
4000 0.0128 0.0146 0.0149
8000 0.0130 0.0133 0.0129
@@ 5 Conclusion
6 Ackonwledgements
We acknowledge financial support from the Jialin Xie Foundation from Institute of high energy physics, Chinese academy of sciences. The authors acknowledge discussions with Haifeng Zhao and Lirong Zheng.